3((s-1)/2)^2<=11 find soluation for s
Answer by mananth(16946)

3((s-1)/2)^2<=11
((s-1)/2)^2<=11/3
(s-1)^2/4 <= 11/3
(s-1)^2<= 44/3
take the square root
s-1 <= +/- (sqrt(44/3))
s<= (+/- sqrt(44/3)) +1